OpenStudy (anonymous):

integral of 6xdx/ sqrt(1-x^2)

OpenStudy (anonymous):

use substitution

OpenStudy (anonymous):

first take out the 6 since its constant, then let u = (1-x^2)

OpenStudy (anonymous):

would it be \[12\sin^{-1}(x) + c \] ?

OpenStudy (anonymous):

or 12xsin −1 (x)+c

OpenStudy (anonymous):

where did u get the sine from?

OpenStudy (anonymous):

it is a rule... where \[\int\limits \frac{ du }{ \sqrt{a^2 -u^2}} + C \]

OpenStudy (anonymous):